OpenStudy (anonymous):

fan and medal Laura is baking a cake. The recipe says that she has to mix 64 grams of chocolate powder to the flour. Laura knows that 1 cup of that particular chocolate powder has a mass of 128 grams. She added two over three of a cup of chocolate powder to the flour. Should Laura add more chocolate powder to make the exact recipe or did she go over and by what amount? She went over by one over three of a cup She needs to add one over three of a cup She needs to add one over six of a cup She went over by one over six of a cup

OpenStudy (studygurl14):

\[\frac{ 1 }{ 128 } = \frac{ 2/3 }{ g }\] \[128(2/3) = g = 85 \frac{ 1 }{ 3 }\] Recipe requires 64 grams. So she went over how much?

OpenStudy (anonymous):

1 over 3 ?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

no, do this: \[(85 \frac{ 1 }{ 3 }) - 64 = ?\]

OpenStudy (anonymous):

1 over 6 ?

OpenStudy (studygurl14):

Okay, ignore the 1/3. Do this: \[85 - 64 = ?\]

OpenStudy (anonymous):

thats not an answer

OpenStudy (studygurl14):

I know. The answer is "?" \[(85 \frac{ 1 }{ 3 }) - 64 = ?\]

OpenStudy (studygurl14):

well, that's the answer in grams. Then you have to convert it back to cups. Understand?

OpenStudy (anonymous):

whats the answer after its converted?

OpenStudy (studygurl14):

1/6, oh you were right! Sorry. I didn't understand what you were doing earlier. So yeah, she went over by 1/6

OpenStudy (anonymous):

ty ty ty ty for your help

OpenStudy (studygurl14):

you're welcome!